OpenStudy (saeed.reichert):

Use signed numbers to solve. A helicopter descended 1200 feet, rose 800 feet, and then descended 450 feet. What was the net gain or loss in altitude? A. lost 2450 ft B. lost 850 ft C. gained 850 ft D. gained 2450 ft

OpenStudy (sloppycanada):

-1200 + 800 = -400 - 450 ft = -850 I would say B, make sense?